# Algebraic non-Hermitian skin effect and unified non-Bloch band theory in arbitrary dimensions

### v1

@@ V The GBZ formula in generic higher dimensional systems
where the tensor $A_{ij}$ represents the superposition coefficients, and $\beta_{x,1}$ and $\beta_{x,2}$ depend on the values of $\rho_{z,i}$ and $\rho_{y,j}$ and satisfies the GBZ condition in Eq. (49). We take the Hamiltonian in Eq. (48) with the lattice size of $L=12$ and set the parameters to $t_{x}=t_{y}=1/2$, $t_{z}=1$, and $t_{xyz}=i/2$. The analytic wavefunction, constructed from the GFS basis, fully aligns with the numerical wavefunction obtained by diagonalizing the Hamiltonian matrix. For verification details, see Appendix B. This successfully verifies the 3D generalization of the 2D GBZ formula. As an illustration, this case features a 3D GFS manifold embedded within a 5D GBZ manifold.
VI Conclusion and Discussion
In summary, this work establishes a unified GBZ formula applicable to any non-Hermitian lattice system across all spatial dimensions, developed using a non-Bloch transfer-matrix approach. We find that in $d$-dimensional non-Hermitian systems, the GBZ dimensionality ranges from $d$ to $12d-$ 1, $d\leq\dim{\text{GBZ}}\leq 2d-1$. Notably, in high dimensional ($d>1$) non-Hermitian systems, the GBZ dimensionality can surpass the space dimension, marking a distinct departure from both Hermitian and 1D non-Hermitian frameworks. This extended GBZ dimensionality provides a platform for new forms of non-Hermitian skin effects in higher dimensions. Within this framework, standing wave conditions can be imposed in one or multiple spatial directions. Using the standing wave recipe, we are enable to precisely construct the open-boundary wavefunctions, which aligns exactly with numerical simulations. This alignment not only provides robust evidence of the GBZ framework’s correctness but also helps elucidate the fundamental characteristics of the high-dimensional skin effect. A key feature we revealed in this work is the algebraic localization in the 2D non-Hermitian skin effect, where the amplitude of skin modes follows a power-law decay with distance from open boundaries.
Contrary to conventional wisdom, which associates the non-Hermitian skin effect with exponential localization at open boundaries and insensitivity to distant local perturbations, the nature of algebraic localization indicates long-range spatial correlations in the higher-dimensional skin effect. Owing to its power-law decay, this universal algebraic skin effect cannot be fully captured by a local imaginary gauge transformation to Bloch waves, underlining its uniqueness and foundational significance in non-Hermitian higher-dimensional systems. The revelation of the long-range correlation of the high-dimensional skin effect is expected to inspire related physical consequences and provide a basis for future research.
Acknowledgements.
We thank C. Fang for valuable discussions. This work was supported in part by the Office of Naval Research (MURI N00014-20-1-2479).
Appendix A The existence of Fermi points ensured by reciprocity
[…]
Due to the reciprocity of the Hamiltonian, the spectral winding number in Eq. (51) satisfies $w(E_{0},k_{y})=-w(E_{0},-k_{y})$. At the high symmetry points ${k^{\ast}_{y}=0,\pi}$, we have $k^{\ast}_{y}=-k^{\ast}_{y}$. Therefore, for a generic OBC eigenvalue $E_{0}$, the spectral winding number satisfies $w(E_{0},k^{\ast}_{y})=-w(E_{0},k^{\ast}_{y})=0$. It means that the PBC spectrum of $h(k_{x},k_{y}^{\ast})$ collapses into arcs in the complex-energy plane Zhang et al. (2020). Similarly, the spectrum of $h(k^{\ast}_{x},k_{y})$ also collapses into arcs at the high symmetry points ${k_{x}^{\ast}=0,\pi}$. Assuming $E_{0}$ is in the inner gap of the PBC spectrum, there is a $k_{y}^{0}$ such that the spectrum of $h(k_{x},k_{y}^{0})$ has a nonzero spectral winding number with respect to $E_{0}$. When we smoothly transition $k_{y}$ from $k_{y}^{0}$ to $k_{y}^{\ast}$ (for example, $k^{\ast}_{y}=0$), the corresponding PBC spectrum of the 1D subsystem of $k_{x}$ must transition smoothly. Therefore, it is impossible to squeeze the spectrum from a loop encircling energy $E_{0}$ (at $k_{y}=k_{y}^{0}$) into an arc (at $k_{y}=k_{y}^{\ast}$) without passing through $E_{0}$. Consequently, the PBC spectrum must sweep through $E_{0}$ at some $k_{y,i}$. These $k_{y,i}$ correspond to the Fermi points of $E_{0}$. Therefore, open-boundary eigenvalue $E_{0}$ must belong to the periodic-boundary spectrum of $h(k_{x},k_{y})$ and the existence of Fermi points is ensured by reciprocity.
Figure 5:  Numerical verification of the 3D GBZ formula is presented using the tight-binding model Hamiltonian in Eq. (B) [or Eq. (48)]. (a) shows the OBC eigenvalues as red dots, with the selected eigenvalue $E_{0}$ marked by a green dot. (b)-(d) illustrate the GFS for the OBC eigenvalue $E_{0}$. (b) shows the transfer matrix eigenvalues $\rho_{z}$, with the selected value $\rho_{z_{0}}$ indicated by a green dot. For $\rho_{z_{0}}$, the 3D Hamiltonian simplifies to a 2D subsystem, and the sub GFS for this subsystem is determined; its projections onto $\rho_{y}$ and $\beta_{x}$ planes are illustrated in (c) and (d), respectively. (e) visualizes the formation of a standing wave component for the 3D OBC eigen wavefunction. The standing wave is formed along the $x$ direction by two equal-amplitude non-Bloch waves, where $|\beta_{x}|=|\widetilde{\beta}_{x}|$. (f) shows the constructed 3D wavefunction of energy $E_{0}$ using the GFS basis, which fully agrees with the numerical wavefunction obtained by diagonalizing the Hamiltonian matrix. The system parameters specified in Eq. (B) are $t_{x}=t_{y}=1/2$, $t_{z}=1$, $t_{xyz}=i/2$, and the system size is $L_{x}=L_{y}=L_{z}=12$.
Appendix B The GBZ formula in three-dimensional systems
Here we provide a detailed numerical verification for GBZ formula of the 3D tight-binding model given by Eq. (48). This section includes two steps: First, we choose the proper planar basis and obtain the GFS/GBZ under this basis. In the most generic cases, $\dim{\text{GFS}}=3$ and $\dim{\text{GBZ}}=5$. Second, for a specific OBC eigenvalue, we can construct the corresponding OBC eigen wavefunction using the GFS basis, which perfectly match the numerical wavefunction obtained by diagonalizing the Hamiltonian matrix.
The real space Hamiltonian for model given by Eq. (48) is written as

### v2

@@ VI The GFS formula for generic higher dimensional lattice systems
where the tensor $A_{ij}$ represents the superposition coefficients, and $\beta_{x,1}$ and $\beta_{x,2}$ depend on the values of $\rho_{z,i}$ and $\rho_{y,j}$ and satisfies the GFS condition in Eq. (49). We take the Hamiltonian in Eq. (48) with the lattice size of $L=12$ and set the parameters to $t_{x}=t_{y}=1/2$, $t_{z}=1$, and $t_{xyz}=i/2$. The analytic wavefunction, constructed from the GFS basis, fully aligns with the numerical wavefunction obtained by diagonalizing the Hamiltonian matrix. See Appendix G for verification details. This verifies the 3D generalization of the 2D GFS formula. As an illustration, this case features a 3D GFS submanifold embedded in a 5D GBZ manifold.
VII The GFS formula in 2D continuous systems
To show the applicability of the GFS formula we developed in lattice models, we now extend it to a two-dimensional continuous system. We demonstrate that the GFS formula can be achieved without the help of the transfer matrix approach.
The strategy is as follows: we first extend the GFS formula to continuous models. Specifically, the GFS curve can be analytically obtained by imposing the standing-wave condition along the $x$ direction. With this choice, the GFS becomes a collection of standing wave bases along the $x$ direction. It’s worth noting that, due to the quantization of the standing wave numbers under OBCs in the $x$ direction, the GFS is inherently an infinite discrete set, even in continuous systems, rather than a continuous set. Using the GFS basis, we can construct the OBC wavefunction in the continuous system with undetermined coefficients. By imposing OBCs in the $y$ direction, we obtain an infinite-dimensional boundary matrix, whose null space determines these coefficients. The existence of a null space (or equivalently, a zero determinant) in this boundary matrix serves as a reliable criterion for identifying the OBC energy spectrum of the continuous system. Finally, we confirm that the numerical wavefunction coincides with the analytic wavefunction constructed from GFS basis.
VII.1 The GFS formula in reciprocal continuous systems
To elaborate on the above approach, we consider the following concrete partial differential equation:
$(-\partial_{x}^{2}-A\,\partial_{y}^{2}-B\,\partial_{x}\partial_{y})\,\psi_{E}(x,y)=E\psi_{E}(x,y),$ (51)
where $A$ and $B$ are generally complex values, thus the differential operator is non-Hermitian. This operator can be expressed in complex-valued momentum space as
$\mathcal{H}(q_{x},q_{y})=q_{x}^{2}+A\,q_{y}^{2}+B\,q_{x}q_{y},$ (52)
where two complex momenta $q_{x}=k_{x}+i\kappa_{x}$ and $q_{y}=k_{y}+i\kappa_{y}$. The linear terms in $q$ in the Hamiltonian are forbidden due to reciprocity: $\mathcal{H}(q_{x},q_{y}){=}\mathcal{H}^{T}(-q_{x},-q_{y})$. Finally, Eq. (52) provides a general expression for reciprocal continuous systems, neglecting higher-order terms. It’s worth noting that although the non-Bloch framework has been applied to 1D continuous systems [102], extending it to two-dimensional non-Hermitian continuous systems remains challenging, due to the lack of a theoretical framework describing the reciprocal non-Hermitian skin effect. In the following, we demonstrate that the GFS formula can be applied to 2D continuous systems. Using the GFS formula, the complex-wave structure of OBC wavefunction and OBC spectrum in the continuous system can be determined.
As the generalization of the GFS formula to continuous system, we first apply the standing-wave condition along the $x$ direction. For a given energy $E_{0}$, there are two branches of complex $q_{x,1}(q_{y},E_{0})$ and $q_{x,2}(q_{y},E_{0})$ by solving $\mathcal{H}(q_{x},q_{y})-E_{0}=0$. The standing-wave condition in the $x$ direction implies that two wave numbers, $k_{x}$ and $k_{x}+\delta k_{x}$, share the same localization length $\kappa_{x}$. Therefore, the $x$-direction standing-wave condition imposes a real constraint on the complex momentum $q_{x}=k_{x}+i\kappa_{x}$:
$\operatorname{Im}q_{x,1}(q_{y},E_{0})=\operatorname{Im}q_{x,2}(q_{y},E_{0}).$ (53)
Within the two real constraints from the characteristic equation $f(q_{x},q_{y},E)=\mathcal{H}(q_{x},q_{y})-E=0$ and one real constraint from the standing-wave condition in Eq. (53), the 4D momentum space ($k_{x},k_{y},\kappa_{x},\kappa_{y}$) reduces to 421=14-2-1=1, yielding a 1D GFS curve.
To obtain the analytic expression of this 1D GFS curve, we utilize the resultant method [32]. The resultant formula works as follows: consider two constraints on three variables $(x_{1},x_{2},x_{3})$, given by $f_{1}(x_{1},x_{2},x_{3})=0$ and $f_{2}(x_{1},x_{2},x_{3})=0$. The common solutions of $f_{1}$ and $f_{2}$ in the $(x_{1},x_{2})$ space are given by the zero locus of their resultant with respect to $x_{3}$: $g(x_{1},x_{2})\equiv\operatorname{Res}_{x_{3}}[f_{1},f_{2}]=0$. Let the complex energy be $E=E_{r}+iE_{i}$, where $E_{r}$ and $E_{i}$ represent the real and imaginary parts of the energy, respectively. Thus, the projection of the GFS on the complex-$q_{x}$ plane can be obtained through the following two resultants:
$g(q_{x},\delta k_{x},E){=}\operatorname{Res}_{q_{y}}[f(q_{x},q_{y},E),f(q_{x}+\delta k_{x},q_{y},E)]$
$R_{x}(k_{x},\kappa_{x},E_{r},E_{i}){=}\operatorname{Res}_{\delta k_{x}}[\operatorname{Re}g,\operatorname{Im}g],$ (54)
where $q_{x},q_{y},E\in\mathbb{C}$ and $\delta,k_{x},\kappa_{x},E_{r},E_{i}\in\mathbb{R}$. Here, $\operatorname{Re}g$ and $\operatorname{Im}g$ denote the real and imaginary parts of the complex function $g$, respectively. Finally, we obtain the real function $R_{x}$. For each given energy $E=E_{r}+iE_{i}$, the equation $R_{x}(k_{x},\kappa_{x},E_{r},E_{i})=0$ determines the projection of the 1D GFS curve in the $k_{x}{-}\kappa_{x}$ plane. As $E$ spans a 2D region in the complex-energy plane, the equation $R_{x}(k_{x},\kappa_{x},E_{r},E_{i})=0$ generates a 3D GBZ manifold. The dimensionality relation follows $\dim{\text{GBZ}}=\dim{\text{GFS}}+\dim{\text{E}}$ given by Eq. (5). Specifically, we set the parameters in Eq. (52) to $A=1+i$ and $B=2$. With these parameters, the continuous model relates to the lattice tight-binding model given by Eq. (39), where the skin modes are localized at bottom and top edges. We select the energy $E_{0}=1/3+i$, the projection of the GFS curve on the $k_{x}{-}\kappa_{x}$ plane is illustrated in Fig. 7(c). Next, the $q_{y}$-projection of the GFS curve for energy $E_{0}$ can be obtained using the resultant method and shown in Fig. 7(d). For each $q_{y}$, there are two $q_{x}$ solutions with the same imaginary part but different real parts, which can form standing waves in the $x$ direction.
Figure 7:  (a) Illustration of Dirichlet boundary conditions (also known as open boundary conditions) in a rectangular geometry with system size $L_{x}=100$ and $L_{y}=120$ for the continuous Hamiltonian given by Eq. (52). The blue solid and dashed lines represent the wavefunction line cuts. (b) The bluish region shows the PBC spectrum, while the red dots represent the first eighty eigenvalues under open boundary conditions. For the selected OBC eigenvalue $E_{0}=1/3+i$, the analytic expression for its generalized Fermi surface (GFS) curves can be obtained using the resultant method. The projections of the GFS curve onto the complex-$q_{x}$ and complex-$q_{y}$ planes are shown in (c) and (d), respectively. The red and blue dots indicate that the GFS curves are parameterized by the $x$-standing wave number $n\in\mathbb{Z}^{+}$. For the numerical OBC wavefunction of $E_{0}$, we extract its components along different line cuts with different values of $b$, as illustrated in (a), and then apply the Fourier transform in Eq. (61). The first ten Fourier coefficients (from $n=1$ to $n=10$) for different line cuts $b$ are displayed in (e). (f) We compare the logarithms of the shift functions $\log[\tilde{A}_{b,2n}/\tilde{A}_{0,2n}]$ obtained from the numerical wavefunction (dots) with the analytical expression (black curves). The excellent agreement between the two confirms the validity of our GFS formula. In (g1) and (g2), we calculate the logarithm of the amplitude of the minimal eigenvalue for the even (odd)-parity boundary matrix, denoted as $\sigma^{+}_{\text{min}}$ ($\sigma^{-}_{\text{min}}$), as the energy moves along the inner (outer) arc trajectory shown in the inset of (b). At angle $\theta_{1}$ ($\theta_{2}$), the trajectory crosses the first OBC eigenvalue of the system, where $\log[\sigma^{+}_{\text{min}}$] ($\log[\sigma^{-}_{\text{min}}]$) reaches its minimum. In (h1) and (h2), we also compute the even (odd)-parity boundary matrix as the energy follows the radial trajectories shown in the inset of (b). At distance $r_{1}$ ($r_{2}$), the trajectory crosses the first(second) OBC eigenvalue, where $\log[\sigma^{+}_{\text{min}}$] ($\log[\sigma^{-}_{\text{min}}]$) also reaches its minimum.
When we impose the zero boundary conditions (or open-boundary condition) on the continuous system with size $0\leq x\leq L_{x}$ and $0\leq y\leq L_{y}$, the standing wave conditions in the $x$ direction quantize the GFS bases into an infinite discrete set, as illustrated in Figs. 7(c)(d). Now we demonstrate the quantization of the GFS bases under the $x$-direction standing-wave condition. The OBC wavefunction can be constructed from the GFS basis as: